OpenStudy (anonymous):

The coordinates of the vertices of quadrilateral ABCD are A(6, 7), B(6, 3), C(3, 2), and D(2, 5). ABCD is reflected across the x-axis to form PQRS. What is the length of diagonal PR? three times the square root of 2 units square root of 34 units square root of 10 units two times the square root of 5 units

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Hero

hero (hero):

Reflect each point across the x axis one at a time.

hero (hero):

\[A \rightarrow P\]\[B \rightarrow Q\]\[C \rightarrow R\]\[D \rightarrow S\]

OpenStudy (anonymous):

(6,-7), (6,-3),(3,-2),(2,-5) @Hero

hero (hero):

P = (6,-7) Q = (6,-3) R = (3,-2) S = (2,-5)

OpenStudy (anonymous):

now how do we get the diagonal? @Hero

hero (hero):

You have to use graph paper and find the length of segments.

OpenStudy (anonymous):

would it be c? @Hero

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

is the answer c?

OpenStudy (phi):

do you know the distance formula to find the distance between 2 points ?

OpenStudy (anonymous):

I tried that but I am super confused and my assignment is due in 5 minutes and this is my only question left! I am desperate for the answer @phi

OpenStudy (phi):

what two points do you use ?

OpenStudy (anonymous):

(6,7) and (6,3)

OpenStudy (anonymous):

@phi

OpenStudy (phi):

those are A and B. (which match up with P and Q) they want diagonal PR

OpenStudy (phi):

and notice 6,7 is 4 steps above 6,3 so it is length 4 which does not match any of your choices... that is a clue you have the wrong points. People have posted up above which points go with P and R

OpenStudy (anonymous):

(6,-7) and (6,-3)

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

sorry I meant to write that the first time you asked

OpenStudy (phi):

those are P and Q you want P and R

OpenStudy (phi):

or A and C

OpenStudy (anonymous):

I got square root of -2 which isnt possible

OpenStudy (anonymous):

i got squareroot of 2

OpenStudy (phi):

show the work

OpenStudy (phi):

first, what 2 points are you using ?